# Ground Every Sentence: Improving Retrieval-Augmented LLMs with Interleaved Reference-Claim Generation

### v1

@@ Abstract
Retrieval-Augmented Generation (RAG) has been widely adopted to enhance Large Language Models (LLMs) in knowledge-intensive tasks. Recently, Attributed Text Generation (ATG) has attracted growing attention, which provides citations to support the model’s responses in RAG, so as to enhance the credibility of LLM-generated content and facilitate verification. Prior methods mainly adopt coarse-grained attributions, linking to passage-level references or providing paragraph-level citations. However, these methods still fall short in verifiability and require certain time costs for fact checking. This paper proposes a fine-grained ATG method called ReClaim (Refer & Claim), which alternates the generation of references and answers step by step. Unlike traditional coarse-grained attribution, ReClaim allows the model to add sentence-level fine-grained citations to each answer sentence in long-form question-answering tasks. Our experiments encompass various training and inference methods and multiple LLMs, verifying the effectiveness of our approach.
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) Lewis et al. (2020) is a technique that integrates information retrieval with natural language generation to enhance the performance and accuracy of large language model (LLMs) responses. However, the RAG system still encounters challenges related to verifiability and credibility. To address these issues, attributed text generation (ATG) Bohnet et al. (2022) has been proposed. ATG aims to improve RAG systems in terms of: 1) Credibility, as explicit citations can reduce hallucinations; 2) Verifiability, making it easier for users to verify the answer.
Previous efforts on ATG mainly focus on passage-level (Thoppilan et al., 2022) or paragraph-level references (Menick et al., 2022; Nakano et al., 2021; Gao et al., 2023b). Although these attribution methods have contributed to improving the verifiability and credibility of model responses, current methods often focus on relatively coarse-grained attributions, which may contain a significant amount of irrelevant information. This increases the time cost for fact-checking.
Figure 1: The task setup for ReClaim. Given a question and reference passages from a large retrieval corpus. The model then generates a text response with fine-grained citations.
In this paper, we propose ReClaim, which generates attributed text with interleaving references and answers for RAG systems, as is shown in Figure 1. This method enables sentence-level fine-grained attributions for the model’s answer in long-form question-answering using the RAG system.
To improve the model’s text generation with citations, we constructed a training dataset, fine-tuned the model, and employed strategies for alternating citation and answer generation. We also added decoding constraints to prevent inconsistencies between citations and source reference passages.
Results of experimental results indicate that our method not only matches but also surpasses existing baseline in some metrics, particularly improving the attribution quality, which refers to the extent to which citations support the answer text. Additionally, our method reduces the length of citations, thereby reducing the effort needed for fact-checking.
@@
    We propose a method called ReClaim, which alternately generates citations and answer sentences, to enable large models to generate answer with citations.
@@
    For ReClaim, we constructed a training dataset and fine-tuned the model using different approaches to improve its attribution capability.
@@
    Through multiple experiments, we demonstrated the effectiveness of our method in enhancing the model’s verifiability and credibility.
@@ 2 Related Work
Our work primarily focuses on the long-form question answering (LFQA) task within the RAG system. Unlike short-form QA Rajpurkar et al. (2016); Joshi et al. (2017), which concentrates on extracting concise facts, LFQA aims to generate comprehensive answers that require a deep understanding of the context and the integration of information from multiple sources. A notable work in this field is the ELI5 (Explain Like I’m Five) dataset Fan et al. (2019), which challenges models to provide straightforward and comprehensible explanations for complex questions. Similarly, the ASQA (Answer Summaries for Questions which are Ambiguous) dataset Stelmakh et al. (2022) requires models to generate abstractive summaries from multiple answer passages, ensuring that the synthesized answer is both coherent and informative.
Generate Text with Citation
Many current works have proposed various methods to generate answer text with citations. These methods differ in their approaches to attribution and the granularity of the citations.
LaMDA Thoppilan et al. (2022) provides attribution for the entire response in the form of URLs pointing to entire documents. WebGPT Nakano et al. (2021) and GopherCite Menick et al. (2022) use reinforcement learning from human preferences to enable LLMs to answer questions while providing snippet citations. ALCE Gao et al. (2023b) goes further by providing one or more input documents as attribution for each sentence in the answer, in a manner similar to cross-referencing.
In addition to the aforementioned methods that add citations directly during answer generation, there are some methods that focus on finding citations afterward. RARR Gao et al. (2023a) uses two steps, Research and Revision, to add attribution to any text post hoc and to edit the original answer based on the found citations.
@@ 3 Method
This study aims to generate text with fine-grained citations by integrating citations with answers in the form of Chain-of-Thought (CoT) Wei et al. (2022). We introduce our approach ReClaim in Section 3.1, and the following sections detail the specific implementation of ReClaim.
@@ 3.1 ReClaim: Interleaving Reference and Claim
Our task can be formally expressed as follows: Given a query q and several reference passages retrieved by the RAG system $\mathcal{D}$, the LLMs is required to generate an output $\mathcal{O}$ = { $\textit{r}_{\textit{1}}$, $\textit{c}_{\textit{1}}$, $\textit{r}_{\textit{2}}$, $\textit{c}_{\textit{2}}$$\textit{r}_{\textit{n}}$, $\textit{c}_{\textit{n}}$, where $\mathcal{O}$ consists of n sentence-level fine-grained references $\textit{r}_{\textit{1}}$, …, $\textit{r}_{\textit{n}}$, which represent the fine-grained citations coming from reference passages and n claims $\textit{c}_{\textit{1}}$, …, $\textit{c}_{\textit{n}}$, which are portions of the model’s response generated based on these references. Each reference $\textit{r}_{\textit{i}}$ corresponds to and serves as substantiation for the corresponding claim $\textit{c}_{\textit{i}}$. All the claims together form the model’s complete answer to the question.
In our work, references are surrounded and denoted by the tags <reference> and </reference>, while claims are demarcated using the tags <claim> and </claim>. Then, they are connected in the form of CoT to produce attributed text response.
During generation, the model needs to alternately generate the reference and claim parts. In the experimental process, it was found that the model encounters several issues when generating references and claims: 1) The references generated by the model are not always completely consistent with the context of the retrieved reference passages; 2) The claims generated by the model do not always attribute well to the corresponding references. Therefore, in the following sections, we will study how to improve the generation of references and claims.
@@ 3.2 Training Dataset Construction
To improve the model’s ability to generate the references and corresponding claims, we constructed a specialized fine-tuning dataset. This dataset was built based on the WebGLM-QA Liu et al. (2023) dataset which consists of 43579 high-quality data samples for the train split. These data samples contain rich reference passages and detailed long-form answer to the question. We performed dataset construction steps and data cleansing processes on this dataset, ultimately resulting in 9433 training samples suitable for our task format.
We adopted a two-stage method to construct our training dataset: 1) Use ChatGPT (gpt-3.5-turbo) OpenAI (2022) to automatically segment the answers in the WebGLM-QA dataset, and identified the corresponding citations in the reference passages based on the segmented clauses; 2) Constructed the attributed answers using the clauses and their corresponding citations in the form of CoT as described in Section 3.1.
To ensure the quality of attribution in the constructed answers, we refined the preliminary training dataset: 1) Segmented the corresponding citation of each clause and performed string matching in the reference passages to filter out mismatched data items where the citation did not align with the original reference passages; 2) Used a natural language inference (NLI) Honovich et al. (2022) model to judge the entailment relationship between the citations (as the premise) and the clauses (as the hypothesis). If there is no implication relationship, it indicates that the citation does not contain sufficient information to support the corresponding clause. Therefore, we need to filter out such data items to ensure a relatively high level of attribution quality.
Through these steps, we ultimately constructed a fine-tuned dataset that ensures text consistency of the references and high attribution quality, providing a foundation for subsequent research.
Samples Average Length
@@
9433 93.6 141.8 282.8
Table 1: Statistics of the training dataset.
@@ 3.3 Unified Generation
[…]
@@ 3.4 Interleaving Generation
During the claim generation stage, since the model has already selected sufficiently granular reference text to follow, which contain the answer information, the full input context is not required. Therefore, the ReClaim w/IG (IG stands for Interleaving Generation) method trains separate models for the generation of the reference parts and the claim parts, and alternates between the two models during answer generation, adjusting the input to each model accordingly. The whole generation process is shown in Figure 2. Specifically, we adopted the following steps to train the models and generate the answer:
Figure 2: The generation process of ReClaim w/IG. Based on the given questions and the reference passages retrieved, the model alternately generates the reference parts and the claim parts in a step-by-step manner. For these two stages of generation, distinct datasets are constructed to train the models, which alternately switches between the models and the input context during inference.
@@
During the generation of the reference parts in the output, the model needs to generate the next reference based on the complete input context and previous output. We define this generation process as $ReferGen=\{\textit{r}_{\textit{i+1}}\ |\ Prompt,\{\textit{r}_{\textit{1}},\textit{c}_{\textit{1}},...,\textit{r}_{\textit{i}},\textit{c}_{\textit{i}}\}\}$, where $\textit{r}_{\textit{i+1}}$ refers to the reference part generated in the next stage, $Prompt$ refers to the complete input context containing instructions, query and reference passages, and $\textit{r}_{\textit{1}},\textit{c}_{\textit{1}},...,\textit{r}_{\textit{i}},\textit{c}_{\textit{i}}$ refer to the previously generated references and claims. As the training of the model for generating the reference parts does not require masking parts of the input context information, we follow the same approach as in the Section 3.3, using the training data constructed in Section 3.2 to fine-tune the model.
@@
During the generation of the claim parts in the output, the model only needs to generate the next claim based on the previous output. We define this generation process as $ClaimGen=\{\textit{c}_{\textit{i+1}}\ |\{\textit{r}_{\textit{1}},\textit{c}_{\textit{1}},...,\textit{r}_{\textit{i}},\textit{c}_{\textit{i}},\textit{r}_{\textit{i+1}}\}\}$, where $\textit{c}_{\textit{i+1}}$ refers to the claim part generated in the next stage, and $\textit{r}_{\textit{1}},\textit{c}_{\textit{1}},...,\textit{r}_{\textit{i}},\textit{c}_{\textit{i}},\textit{r}_{\textit{i+1}}$ refer to the previously generated references and claims. We use the training data constructed in Section 3.2 and split it into the data format that conforms to our described generation process ClaimGen. This formatted dataset is then used to fine-tune the model.
After generating the two models mentioned above, we can alternate between these models during the inference stage while simultaneously adjusting the model input according to the defined generation process. This allows us to separately generate the reference and claim parts, ultimately producing attributed text output.
@@ 3.5 Decoding Constraints
To ensure the generated reference parts align with the source reference passages, we impose decoding constraints through the following three steps:
@@
We segment the reference passages into individual sentences. Then, we use the model tokenizer to encode these sentences into vectors. Each vector representation of a sentence serves as the smallest unit for generating the reference parts.
@@
During the model inference stage for generating reference parts, we select the token with the highest generation probability that satisfies the current prefix tree path as the next output token. This process continues until reaching a leaf node. Upon reaching a leaf node, the model either select another prefix tree path for output or begin the claim generation. This approach allows us to select a complete and consistent sentence from the reference passages as part of our current reference each time.
@@ 4 Experimental Protocol
In this section, we employ the GPT models and several open-source models to validate the effectiveness of our method across multiple evaluation dimensions. We conduct a comprehensive analysis by assessing the performance of our approach on various metrics.
@@ 4.1 Evaluation Datasets
Our method primarily targets long-form question answering task within the RAG system, aiming to generate attributed answer. Therefore, we selected ASQA Stelmakh et al. (2022) and ELI5 Fan et al. (2019) as our main benchmark datasets. By testing on these datasets, we can comprehensively assess the effectiveness of our method in generating accurate, coherent, and well-attributed long-form answers.

### v2

@@ Abstract
Retrieval-Augmented Generation (RAG) has been widely adopted to enhance Large Language Models (LLMs) in knowledge-intensive tasks. To enhance credibility and verifiability in RAG systems, Attributed Text Generation (ATG) is proposed, which provides citations to retrieval knowledge in LLM-generated responses. Prior methods mainly adopt coarse-grained attributions, with passage-level or paragraph-level references or citations, which fall short in verifiability. This paper proposes ReClaim (Refer & Claim), a fine-grained ATG method that alternates the generation of references and answers step by step. Different from previous coarse-grained attribution, ReClaim provides sentence-level citations in long-form question-answering tasks. With extensive experiments, we verify the effectiveness of ReClaim in extensive settings, achieving a citation accuracy rate of 90%. [^1] Code and datasets are public at: https://github.com/pdxthree/ReClaim
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) Lewis et al. (2020) is a technique that integrates information retrieval with natural language generation to enhance the performance of large language model (LLMs) responses. However, the RAG system still encounters challenges related to verifiability and credibility. To address these issues, attributed text generation (ATG) Bohnet et al. (2022) has been proposed. ATG aims to improve RAG systems in terms of: 1) Credibility, as explicit citations can reduce hallucinations; 2) Verifiability, making it easier for users to verify the answer.
Previous efforts on ATG mainly focus on passage-level (Thoppilan et al., 2022) or paragraph-level references (Menick et al., 2022; Nakano et al., 2021; Gao et al., 2023b). Although these attribution methods have contributed to improving the verifiability and credibility of model responses, current methods often focus on relatively coarse-grained attributions, which may contain a significant amount of irrelevant information. This increases the time required for fact-checking.
Figure 1: The task setup for ReClaim. Given question and reference passages from a large corpus. The LLM then generates a response with fine-grained citations. For detailed examples, see Table LABEL:table:testdatacase.
In this paper, we propose ReClaim, which generates attributed text with interleaving references and answers for RAG systems, as is shown in Figure 1. This method enables sentence-level fine-grained attributions in long-form question-answering using the RAG system.
To enhance the LLM’s performance in attributed text generation, we developed a training dataset and fine-tuned the LLM to facilitate sentence-level citation selection from given reference passages and subsequent answer generation. We implemented an alternating strategy between citation generation and answer text generation. We apply constrained decoding during LLM inference by encoding reference passages into a token-level prefix tree. This restricts the LLM to generate citations only along the tree’s paths, ensuring alignment with the reference passages and avoiding inconsistencies.
The results of experiments demonstrate that ReClaim outperforms existing baselines. ReClaim significantly improves the citation quality, enabling the citations to better support the answer. Furthermore, ReClaim greatly reduces the verbosity of citations, thereby easing the fact-checking process.
@@
    We propose a method called ReClaim, which alternately generates citations and answer sentences, enabling LLM to produce answers with sentence-level citations, thus enhancing the LLM’s verifiability and credibility.
@@
    To enhance LLMs in sentence-level citation generation, we construct a dataset based on WebGLM-QA Liu et al. (2023) and ELI5 Fan et al. (2019) dataset. Then, we fine-tune Llama3-8B-Instruct Dubey et al. (2024) models for reference and claim generation respectively, achieving better citation quality compared to the baseline method with ChatGPT OpenAI (2022).
@@
    Through extensive experiments, we demonstrate the effectiveness of our method in enhancing the LLM’s verifiability and credibility, achieving performance comparable to much bigger models like ChatGPT.
@@ 2 Related Work
Our work primarily focuses on the long-form question answering (LFQA) task within the RAG system. Unlike short-form QA Rajpurkar et al. (2016); Joshi et al. (2017), which concentrates on extracting concise facts, LFQA generates comprehensive answers that require deep contextual understanding and information integration from multiple sources Fan et al. (2019); Stelmakh et al. (2022); Malaviya et al. (2024).
Attributed Text Generation
Many current works propose various methods for generating answer text with citations, differing in their approaches to attribution and citation granularity.
LaMDA Thoppilan et al. (2022) provides attribution for the entire response in the form of URLs pointing to entire documents. WebGPT Nakano et al. (2021) and GopherCite Menick et al. (2022) use reinforcement learning from human preferences to enable LLMs to answer questions while providing snippet citations. ALCE Gao et al. (2023b) goes further by providing one or more input documents as attribution for each sentence in the answer, in a manner similar to cross-referencing. Additionally, some work has focused on fine-tuning models to improve the generation of attributed answer text Huang et al. (2024a); Asai et al. (2023); Huang et al. (2024c).
In addition to the aforementioned methods that add citations directly during answer generation, there are some works that focus on finding citations afterward Gao et al. (2023a). Some research further achieves better attribution performance through multiple retrievals and validations Li et al. (2024); Sun et al. (2023).
@@ 3 Method
We introduce ReClaim, which aims to generate text with sentence-level citations. Specifically, ReClaim generates citations with answers alternatively, in a step-by-step manner. We first introduce the overview of ReClaim in Section 3.1, and then detail the specific implementation in Section 3.2 to 3.5.
Figure 2: The generation process of ReClaim w/IG. Based on the given questions and the reference passages retrieved, the LLM alternately generates the reference parts and the claim parts in a step-by-step manner. For these two stages of generation, distinct datasets are constructed to train the base model, which alternately switches between the fine-tuned models and the input context during inference.
@@ 3.1 ReClaim: Interleaving Reference and Claim
Our task can be formally expressed as follows: Given a query q and several reference passages retrieved by the RAG system $\mathcal{D}$, the LLM is required to generate an output $\mathcal{O}$ = { $\textit{r}_{\textit{1}}$, $\textit{c}_{\textit{1}}$, $\textit{r}_{\textit{2}}$, $\textit{c}_{\textit{2}}$$\textit{r}_{\textit{n}}$, $\textit{c}_{\textit{n}}$, where $\mathcal{O}$ consists of n sentence-level fine-grained references $\textit{r}_{\textit{1}}$, …, $\textit{r}_{\textit{n}}$, which represent the fine-grained citations coming from reference passages (provided as text, not just source numbers) and n claims $\textit{c}_{\textit{1}}$, …, $\textit{c}_{\textit{n}}$, which are portions of the LLM’s response generated based on these references. Each reference $\textit{r}_{\textit{i}}$ serves as a substantiation of claim $\textit{c}_{\textit{i}}$, and together, all $\textit{c}_{\textit{i}}$ form the LLM’s complete answer to the question.
During generation, the LLM alternates between generating references and claims. However, experiments revealed that the LLM faces several issues: 1) The generated references are not always consistent with the retrieved passages; 2) The generated claims do not always attribute well to the corresponding references. Therefore, in the following sections, we study how to improve the generation of references and claims.
@@ 3.2 Training Dataset Construction
To improve the LLM’s ability to choose the references and generate corresponding claims, we construct a specialized fine-tuning dataset based on the WebGLM-QA Liu et al. (2023) and ELI5 Fan et al. (2019) datasets.
The training data input includes an instruction, a query, and reference passages. The output consists of multiple reference-claim pairs, with the reference being a sentence-level citation and the claim being a part of the answer. The reference and claim components are designed to train the LLM’s ability to select accurate and relevant fine-grained citations from the complete reference passages, and to generate fluent, coherent answer texts that are faithful to the selected citations.
The steps of our dataset construction are detailed as follows:
Reference Passages Retrieval
In the WebGLM-QA Dataset, the reference passages are fine-grained texts closely aligned with the query. However, some research has shown that irrelevant reference texts significantly impact the quality of LLM-generated answers Wu et al. (2024). To enhance the LLM’s performance against irrelevant contexts, we sample a portion of the constructed WebGLM-QA training dataset, retrieve the top-100 passages using BM25, and calculate the similarity between the query and passages with a Reranker model Chen et al. (2024). Passages with high BM25 ranking but low re-rank score are selected as irrelevant text and added to the training data.
For the ELI5 dataset, we use BM25 to retrieve the top-100 reference passages, which included some irrelevant noise to simulate real-world retrieval conditions. Additionally, we use the Reranker to re-rank the top 100 passages retrieved from a subset of the selected ELI5 dataset, producing a refined re-ranked dataset.
Model Answer Generation
For the WebGLM-QA dataset, we directly use the original model responses. For the ELI5 datasets, due to significant information discrepancies between the human answers and the retrieved passages, we provide the top-5 retrieved passages to the LLM (Llama-3.1-405B-Instruct Dubey et al. (2024)) to generate long-form answers.
Multi-Stage Citation Search
To select sentence-level fine-grained citations for the answers generated by the LLM, we employ a multi-stage citation search approach that moves from coarse to fine.
We first enable the Llama-3.1-405B-Instruct model to automatically segment the text into clauses. For each clause, the model identifies the minimal set of citations from the reference passages that sufficiently support it.
Then, we use NLI (Natural Language Inference) model Honovich et al. (2022) to evaluate the entailment relationship between the chosen citations (as the premise) and the corresponding clause (as the hypothesis). To enhance filtering precision, we set a threshold ($\theta=0.8$) to discard cases where the entailment probability between the chosen citation and the clause falls below the threshold.
Dataset Samples Average Length
@@
WebGLM-QA 4582 98.53 154.62 326.57
WebGLM-QA Extend 2605 83.85 114.1 396.16
ELI5 Default 3383 91.33 132.34 545.01
ELI5 Rerank 2653 107.54 158.52 542.02
Table 1: Statistics of the training dataset. For more details, please refer to the appendix B.
@@ 3.3 Unified Generation
The ReClaim ${}_{\texttt{Unified}}$ method uses a simple fine-tuning and inference approach. It first performs instruction fine-tuning on the LLM using the dataset constructed in Section 3.2. Then, it uses the fine-tuned LLM to perform one-step generation. Based on the given query and reference passages, it directly outputs the attributed answer. This generation process can be described as: $UnifiedGen=\{\{\textit{r}_{\textit{1}},\textit{c}_{\textit{1}},...,\textit{r}_{\textit{i}},\textit{c}_{\textit{i}}\}\ |\ Query,Passages\}$.
@@ 3.4 Interleaving Generation
During the claim generation stage, since the LLM has already selected sufficiently granular reference text to follow, which contain the answer information, the full input context is not required. Therefore, the ReClaim w/IG method trains separate LLMs for the generation of the reference parts and the claim parts, and alternates between the two LLMs during answer generation, adjusting the input to each LLM accordingly (IG represents the interleaving use of two fine-tuned LLMs for the iterative generation of references and claims).
The entire generation process, as illustrated in Figure 2, involves the following steps to train the LLMs and generate the answer.
@@
During the generation of reference parts, the LLM needs to generate the next reference based on the complete input context and previous output. We define this generation process as $ReferGen=\{\textit{r}_{\textit{i+1}}\ |\ Prompt,\{\textit{r}_{\textit{1}},\textit{c}_{\textit{1}},...,\textit{r}_{\textit{i}},\textit{c}_{\textit{i}}\}\}$, where $\textit{r}_{\textit{i+1}}$ refers to the reference part generated in the next stage, $Prompt$ refers to the complete input context containing instructions, query and reference passages, and $\textit{r}_{\textit{1}},\textit{c}_{\textit{1}},...,\textit{r}_{\textit{i}},\textit{c}_{\textit{i}}$ refer to the previously generated references and claims. As the training of the LLM for generating the reference parts does not require masking parts of the input context information, we follow the same approach as in the Section 3.3 to fine-tune the ReferModel.
@@
During the generation of claim parts, the LLM only needs to generate the next claim based on the previous output. We define this generation process as $ClaimGen=\{\textit{c}_{\textit{i+1}}\ |\{\textit{r}_{\textit{1}},\textit{c}_{\textit{1}},...,\textit{r}_{\textit{i}},\textit{c}_{\textit{i}},\textit{r}_{\textit{i+1}}\}\}$, where $\textit{c}_{\textit{i+1}}$ refers to the claim part generated in the next stage, and $\textit{r}_{\textit{1}},\textit{c}_{\textit{1}},...,\textit{r}_{\textit{i}},\textit{c}_{\textit{i}},\textit{r}_{\textit{i+1}}$ refer to the previously generated references and claims. We utilize the training dataset from Section 3.2, formatting it to align with our ClaimGen generation process, and then use this formatted dataset to fine-tune the ClaimModel.
@@ 3.5 Decoding Constraints
To ensure the generated reference parts align with the retrieved reference passages, we apply decoding constraints through the following three steps:
@@
We segment the reference passages into individual sentences. Then, we use the LLM tokenizer to encode these sentences into vectors. Each vector representation of a sentence serves as the smallest unit for generating the reference parts.
@@
During the inference stage for generating reference parts, we select the token with the highest generation probability that satisfies the current prefix tree path as the next output token. This process continues until reaching a leaf node. Upon reaching a leaf node, the LLM either select another prefix tree path for output or begin the claim generation. This approach allows us to select a complete and consistent sentence from the reference passages as part of our current reference each time.
@@ 4 Experimental Protocol
In this section, we employ the GPT models and several open-source LLMs to validate the effectiveness of our method across multiple evaluation dimensions. We conduct a comprehensive analysis by assessing the performance of our approach on various metrics.
@@ 4.1 Evaluation Datasets
ASQA
We evaluate the 948 samples from the ASQA dataset Stelmakh et al. (2022), selected by ALCE, and use the five oracle paragraphs provided by ALCE as reference passages, which are chosen from the top 100 retrieved passages representing the gold passages.
ELI5
We evaluate the 1,000 samples selected from ELI5 dataset Fan et al. (2019) by ALCE and use the five oracle passages as reference passages.
EXPERTQA
To test the LLM’s generalization ability under the standard RAG process, we select 1,000 samples from the EXPRTQA Malaviya et al. (2024) dataset. We follow the standard RAG procedure: using BM25 to retrieve the top-100 passages, then re-ranking them to select the top 5 passages.